### v1

# On the edge densities of convex mosaics

Abstract.
In this paper we investigate the problem of finding the minimum of edge density in families of convex mosaics with unit volume cells in $3$-dimensional Euclidean space. We solve this problem for two families. In particular, we prove that a translative convex mosaic with unit volume cells has minimal edge density if and only if it is a face-to-face mosaic generated by unit cubes. We also investigate decomposable mosaics, defined as mosaics whose cells are Minkowski sums of cells of two lower dimensional convex mosaics, and show that the edge density of every normal decomposable mosaic with unit volume cells is at least as large as that of a face-to-face mosaic with regular triangle-based right prisms as cells. In addition, we use our method to determine the minimal total edge length of a $3$-dimensional unit volume parallelohedron.
Key words and phrases:
Honeycomb Conjecture, Kelvin’s conjecture, edge density, translative mosaic, decomposable mosaic, parallelohedron, total edge length, Melzak’s conjecture
2020 Mathematics Subject Classification:
52C22, 52A40, 52A38
Partially supported by the National Research, Development and Innovation Office, NKFI, K-147544 grant, and BME Water Sciences & Disaster Prevention TKP2020 Institution Excellence Subprogram, grant no. TKP2020 BME-IKA-VIZ
The aim of this paper is to investigate certain properties of mosaics. A mosaic or tiling of the Euclidean $n$-space $\mathbb{R}^{n}$ is a countable family of compact sets $\mathcal{M}$, called cells, with the property that $\bigcup\mathcal{M}=\mathbb{R}^{n}$, and the interiors of any two distinct cells are disjoint. A mosaic is convex, if every cell in it is convex; it is known that in this case every cell is a convex polytope. If, for a mosaic $\mathcal{M}$, there are universal constants $0<r<R$ such that the inradius of every cell of $\mathcal{M}$ is at least $r$, and its circumradius is at most $R$, the mosaic is called normal. In this paper we deal only with normal, convex mosaics. A convex mosaic is called face-to-face (in the plane also edge-to-edge) if every facet of every cell is a facet of exactly one more cell. Furthermore, a convex mosaic is translative (resp. congruent), if every cell is a translate (resp. a congruent copy) of a given convex polytope. Clearly, every congruent and translative convex mosaic is normal. A cell of a translative convex mosaic in $\mathbb{R}^{n}$ is called an $n$-dimensional parallelohedron.
One of the classic problems of geometry, regarding planar mosaics, is the Honeycomb Conjecture, stating that in a decomposition of the Euclidean plane into cells of equal area, the regular hexagonal grid has minimal edge density. This conjecture first appeared in Roman time in a book of Varro about agriculture [26]. The problem has been the centre of research throughout the second half of the 20th century [7, 19]; it was solved by Fejes Tóth for normal, convex mosaics, while the most general version is due to Hales [12], who dropped the condition of convexity.
A famous analogous problem for mosaics in the $3$-dimensional Euclidean space $\mathbb{R}^{3}$ was proposed by Lord Kelvin [14] in 1887; he conjectured that in a tiling of space with cells of unit volume, minimal surface density is attained by a mosaic composed of slightly modified truncated octahedra. Even though this conjecture was disproved by Weaire and Phelan in 1994 [28], who discovered a tiling of space with two slighly curved ‘polyhedra’ of equal volume and with less surface density than in Lord Kelvin’s mosaic, the original problem of finding the mosaics with equal volume cells and minimal surface density has been extensively studied (see, e.g. [15, 9, 21]). On the other hand, in the author’s knowledge, there is no subfamily of mosaics for which Kelvin’s problem is solved. Here, in particular, we should recall a yet unsolved conjecture of Bezdek [3] from 2006, restated in [18] in 2022, that states that among translative, convex mosaics with unit volume cells, the ones generated by regular truncated octahedra have minimal surface density.
We intend to investigate a similar problem, namely we search for convex mosaics with unit volume cells and minimal edge density. In our paper, we denote the edge density of a mosaic $\mathcal{M}$ by $\rho_{1}(\mathcal{M})$ (for a formal definition of edge density, see Section 1), and the family of translative convex mosaics of $\mathbb{R}^{3}$ with unit volume cells by $\mathcal{F}_{tr}$, and the family of unit volume parallelohedra by $\mathcal{P}_{tr}$.
Our first result is as follows.
Theorem 1.
For any $\mathcal{M}\in\mathcal{F}_{tr}$,
$\rho_{1}(\mathcal{M})\geq 3,$
with equality if and only if $\mathcal{M}$ is a face-to-face mosaic with unit cubes as cells.
To prove Theorem 2, we prove a more general theorem. To state this theorem, we recall a few properties of $3$-dimensional parallelohedra. First, we note that every $3$-dimensonal parallelohedron $P$ is the Minkowski sum of at most $6$ segments. Furthermore, every edge of $P$ is a translate of a generating segment of $P$, and for every generating segment $S$ of $P$ there are $6$ or $4$ edges of $P$ that are translates of $S$. In these cases we say that $S$ generates a $6$-belt or a $4$-belt, respectively.
If $P=\sum_{i=1}^{k}S_{i}$ is a $3$-dimensional parallelohedron generated by the segments $S_{i}$, and $m=(\alpha_{6},\alpha_{4})$ has positive coordinates, we set
$w_{i}=\left\{\begin{array}[]{l}\alpha_{6}\hbox{ if }S_{i}\hbox{ generates a }6-\hbox{belt;}\\ \alpha_{4}\hbox{ if }S_{i}\hbox{ generates a }4-\hbox{belt.}\end{array}\right.$
Furthermore, we define the function $w_{m}(P)=\sum_{i=1}^{k}w_{i}L(S_{i})$, where $L(S_{i})$ denotes the length of $S_{i}$. We prove Theorem 2, in which $P_{cu}$ denotes a unit cube, $P_{oc}$ denotes a unit volume regular truncated octahedron, and $P_{pr}$ denotes a regular hexagon-based right prism, where the edge length of the hexagon is $\frac{2^{2/3}\alpha_{6}^{1/3}}{3^{5/6}\alpha_{4}^{1/3}}$, and the length of the lateral edges of $P_{pr}$ is $\frac{3^{1/6}\alpha_{4}^{2/3}}{2^{1/3}\alpha_{6}^{2/3}}$. We note that the volume of $P_{pr}$ is also one.
Theorem 2.
Let $m=(\alpha_{6},\alpha_{4})$ with $\alpha_{4},\alpha_{6}>0$. Then $w_{M}(\cdot)$ attains its minimum on $\mathcal{P}_{tr}$. Furthermore, if $P\in\mathcal{P}_{tr}$ minimizes $w_{m}(\cdot)$ on $\mathcal{P}_{tr}$, then the following holds.
  • (a)
    If $0<\alpha_{4}<\frac{\sqrt{3}}{2}\alpha_{6}$, then $P$ is congruent to $P_{cu}$.
  • (b)
    If $\alpha_{4}=\frac{\sqrt{3}}{2}\alpha_{6}(\approx 0.866025\alpha_{6})$, then $P$ is congruent to $P_{cu}$ or $P_{pr}$.
  • (c)
    If $\frac{\sqrt{3}}{2}\alpha_{6}<\alpha_{4}<\sqrt[4]{\frac{2}{3}}\alpha_{6}$, then $P$ is congruent to $P_{pr}$.
  • (d)
    If $\alpha_{4}=\sqrt[4]{\frac{2}{3}}\alpha_{6}(\approx 0.903202\alpha_{6})$, then $P$ is congruent to $P_{pr}$ or $P_{oc}$.
  • (e)
    If $4]{\frac{2}{3}}\alpha_{6}<\alpha_{4}$, then $P$ is congruent to $P_{oc}$.
We note that if $m=\left(\frac{1}{2},\frac{1}{2}\right)$, then $w_{m}(P)$ coincides with the mean width of $P$ for any $P\in\mathcal{P}_{tr}$ [18], and thus, Theorem 2 is a generalization of the main result of [18], stating that among $3$-dimensional unit volume parallelohedra, regular truncated octahedra have minimal mean width.
In the paper we also deal with another kind of mosaics.
Definition 1.
A convex mosaic $\mathcal{M}$ of the Euclidean $n$-space $\mathbb{R}^{n}$ is said to be decomposable, if there are linear subspaces $L_{1},L_{2}$ of dimensions $k$ and $n-k$, respectively, with $0<k<n$ and $L_{1}\cap L_{2}=\{o\}$, and convex mosaics $\mathcal{M}_{1}$ of $L_{1}$ and $\mathcal{M}_{2}$ of $L_{2}$ such that a set $C$ is a cell of $\mathcal{M}$ if and only if $C=C_{1}+C_{2}$ for some cells $C_{i}$ of $\mathcal{M}_{i}$, where $i=1,2$.
In the paper, we denote the family of normal, decomposable mosaics of $\mathbb{R}^{3}$ with unit volume cells by $\mathcal{F}_{dc}$, and by $P_{pr}$ we denote a right prism whose base is a regular triangle of edge length $\frac{1}{\sqrt{3}}$, and whose lateral edges have length $\frac{2}{\sqrt{3}}$.
Theorem 3.
For any $\mathcal{M}\in\mathcal{F}_{dc}$,
$\rho_{1}(\mathcal{M})\geq\frac{3\sqrt{3}}{2}.$
Here, equality is attained e.g. if $\mathcal{M}$ is a face-to-face mosaic the cells of which are congruent to $P_{pr}$.
We note that a simplified version of the proof of Theorem 3 yields a proof of the Honeycomb Conjecture for normal, convex mosaics. Our results are also related to a conjecture of Melzak [16] made in 1965, claiming that among convex polyhedra of unit volume, regular triangle based prisms, with equal-length edges, have minimal total edge length. This conjecture is still very far from settled; for the solution of this problem in smaller families of convex polyhedra and for bounds on this quantity, see e.g. [1, 2, 4, 8, 17].
Corollary 1 is an immediate consequence of Theorem 2, and Corollary 2 can be proved by a much simplified variant of the proof of Theorem 3. Here, by $L(P)$ we denote the total edge length of the convex polyhedron $P$.
Corollary 1.
For any $3$-dimensional parallelohedron $P$ of unit volume, we have
$L(P)\geq 12,$
with equality if and only if $P$ is a unit cube.
Corollary 2.
For any prism $P$ of unit volume, we have
$L(P)\geq 2^{2/3}3^{11/6}\approx 11.8962,$
with equality if and only if $P$ is a regular triangle based right prism, and every edge of $P$ is of length $\frac{2^{2/3}}{3^{1/6}}$.
We make the following conjecture.
Conjecture 1.
For any normal, convex mosaic $\mathcal{M}$ with unit volume cells, we have
$\rho_{1}(\mathcal{M})\geq\frac{3\sqrt{3}}{2}.$
The structure of our paper is as follows. In Section 1 we introduce the concepts used in the paper and collect the necessary tools to prove our results. In Section 2 we show how Theorem 2 implies Theorem 1. Finally, in Sections 3 and 4 we prove Theorems 2 and 3, respectively.
1. Preliminaries
Let $\mathcal{M}$ be a convex mosaic in $\mathbb{R}^{3}$. We define the edge density of $\mathcal{M}$ as the limit
$\rho_{1}(\mathcal{M})=\lim_{R\to\infty}\frac{L(\operatorname{skel}_{1}(\mathcal{M})\cap(R\mathbf{B}^{3}))}{\operatorname{vol}(R\mathbf{B}^{3})},$
if it exists, where $\mathbf{B}^{3}$ is the closed unit ball centered at the origin $o$, the $1$-skeleton $\operatorname{skel}_{1}(\mathcal{M})$ of $\mathcal{M}$ is defined as the union of all edges of all cells of $\mathcal{M}$, $L(\cdot)$ denotes length, and $\operatorname{vol}(\cdot)$ denotes volume. We note that if $\mathcal{M}$ is normal, then $\rho_{1}(\mathcal{M})$ exists.
It is shown in [20, 27] that an $n$-dimensional convex polytope $P$ is a parallelohedron if and only if
  • (i)
    $P$ and all its facets are centrally symmetric, and
  • (ii)
    the projection of $P$ along any of its $(n-2)$-dimensional faces is a parallelogram or a centrally symmetric hexagon.
It is worth noting that any affine image of a parallelohedron is also a parallelohedron, and by [5], for any $n$-dimensional parallelohedron $P$ there is a face-to-face tiling of $\mathbb{R}^{n}$ with translates of $P$ as cells.
The combinatorial classes of $3$-dimensional parallelohedra were described a long time ago [6]. More specifically, any $3$-dimensional parallelohedron is combinatorially isomorphic to one of the following (see Figure 1):
  1. (1)
    a cube,
  2. (2)
    a hexagonal prism,
  3. (3)
    Kepler’s rhombic dodecahedron (which we call a regular rhombic dodecahedron),
  4. (4)
    an elongated rhombic dodecahedron,
  5. (5)
    a regular truncated octahedron.
We call a parallelohedron combinatorially isomorphic to the polyhedron in (i) a type (i) parallelohedron.
Figure 1. The five combinatorial types of $3$-dimensional parallelohedra. The type (5) parallelohedron in the picture is the regular truncated octahedron generated by the six segments connecting the midpoints of opposite edges of a cube. The type (3) polyhedron is the regular rhombic dodecahedron generated by the four diagonals of the same cube. The rest of the polyhedra in the picture are obtained by removing some generating segments from the type (5) polyhedron.
Every $3$-dimensional parallelohedron is a zonotope; in particular, a type (1)-(5) parallelohedron is the Minkowski sum of $3,4,4,5,6$ segments, respectively. A typical parallelohedron is of type (5), as every other parallelohedron can be obtained by removing some of the generating vectors of a type (5) parallelohedron, or setting their lengths to zero.
By well-known properties of zonotopes (see e.g. [13, 18, 24]), every edge of a $3$-dimensional parallelohedron $P$ is a translate of one of the segments generating $P$. Furthermore, for any generating segment $S$, the faces $F_{i}$ of $P$ that contain a translate of $S$ can be written in a cyclic sequence $F_{1},F_{2},\ldots,F_{s}=F_{0}$ such that for each $i$, $F_{i}\cap F_{i+1}$ is a translate of $S$; the union of these faces is called the zone of $S$. By property (ii) of parallelohedra, for any such zone the number of elements is $4$ or $6$, implying that for any generating segment $S$ of $P$ there are $4$ or $6$ translates of $S$ among the edges of $P$. In these cases we say that $S$ generates a $4$-belt or a $6$-belt, respectively. For $i=1,2,3,4,5$, a type (i) parallelohedron contains $3,3,0,1,0$ $4$-belts, and $0,1,4,4,6$ $6$-belts, respectively.
To prove Theorem 2 we adapt the representation of $3$-dimensional parallelohedra introduced in [18]. More specifically, if $P$ is a $3$-dimensional parallelohedron, then, up to translation, there are vectors $v_{1},v_{2},v_{3},v_{4}\in\mathbb{R}^{3}$ such that
(1) $P=\sum_{1\leq i<j\leq 4}[o,\beta_{ij}(v_{i}\times v_{j})]$
for some non-negative real numbers $\beta_{ij}$. Here, if $P$ is a type (5) parallelohedron, the vectors $v_{i}$ are normal vectors of the pairs of hexagon faces. As it was shown in [18], we may assume that any three of the $v_{i}$ are linearly independent, and satisfy $v_{1}+v_{2}+v_{3}+v_{4}=o$. We may also assume that $V_{123}=1$, where $V_{ijk}$ denotes the value of the determinant with columns $v_{i},v_{j}v_{k}$. Then, setting $T=\operatorname{conv}\{v_{1},v_{2},v_{3},v_{4}\}$, $T$ is centered, i.e. its center of mass is $o$, implying that for any pairwise different indices $i,j,k$, we have $|V_{ijk}|=1$. The sign of $V_{ijk}$ can be determined from our choice that $V_{123}=1$, the properties of determinants, and from the identity that for any $\{i,j,s,t\}=\{1,2,3,4\}$, we have $V_{ijs}=-V_{ijt}$.
Every $3$-dimensional parallelohedron can be obtained from a type (5) parallelohedron by setting some of the $\beta_{ij}$s to be zero. In particular, $P$ is of type (5), or (4), or (3), or (2), or (1) if and only if
  • $\beta_{ij}>0$ for all $i\neq j$, or
  • exactly one of the $\beta_{ij}$s is zero, or
  • exactly two of the $\beta_{ij}$s is zero: $\beta_{i_{1}j_{1}}=\beta_{i_{2},j_{2}}=0$, and they satisfy $\{i_{1},j_{1}\}\cap\{i_{2},j_{2}\}=\emptyset$, or
  • exactly two of the $\beta_{ij}$s is zero: $\beta_{i_{1}j_{1}}=\beta_{i_{2},j_{2}}=0$, and they satisfy $\{i_{1},j_{1}\}\cap\{i_{2},j_{2}\}\neq\emptyset$, or
  • exactly three of the $\beta_{ij}$s are zero, and there is no $s\in\{1,2,3,4\}$ such that the indices of all nonzero $\beta_{ij}$s contain $s$, respectively.
In the remaining cases $P$ is planar.
The volume of a parallelohedron $P$ in the above representation is independent of the $v_{i}$. More specifically,
(2) $\operatorname{vol}_{3}(P)=f(\beta_{12},\beta_{13},\beta_{14},\beta_{23},\beta_{24},\beta_{34}),$
where the function $f:\mathbb{R}^{6}\to\mathbb{R}$ is defined as
(3) $f(\tau_{12},\tau_{13},\tau_{14},\tau_{23},\tau_{24},\tau_{34})=\tau_{12}\tau_{13}\tau_{23}+\tau_{12}\tau_{14}\tau_{24}+\tau_{13}\tau_{14}\tau_{34}+\tau_{23}\tau_{24}\tau_{34}+\\ +(\tau_{12}+\tau_{34})(\tau_{13}\tau_{24}+\tau_{14}\tau_{23})+(\tau_{13}+\tau_{24})(\tau_{12}\tau_{34}+\tau_{14}\tau_{23})+(\tau_{14}+\tau_{23})(\tau_{12}\tau_{34}+\tau_{13}\tau_{24}).$
We remark that $f$ is not the third elementary symmetric function on six variables, since it has only $16<20=\binom{6}{3}$ members.
Lemma 1 is proved in [18].
Lemma 1.
Let $T=\operatorname{conv}\{p_{1},p_{2},p_{3},p_{4}\}$ be an arbitrary centered tetrahedron with volume $V>0$ where the vertices are labelled in such a way that the determinant with columns $p_{1},p_{2},p_{3}$ is positive. For any $\{i,j,s,t\}=\{1,2,3,4\}$, let $\gamma_{ij}=-\langle p_{s},p_{t}\rangle$ and $\zeta_{ij}=\gamma_{ij}|p_{i}\times p_{j}|^{2}$. Then
  1. (1)
    $f(\gamma_{12},\gamma_{13},\gamma_{14},\gamma_{23},\gamma_{24},\gamma_{34})=\frac{9}{4}V^{2}$,
  2. (2)
    $\sum_{1\leq i<j\leq 4}\zeta_{ij}=\frac{27}{4}V^{2}$,
where $f$ is the function defined in (3).
To prove our result, we need the following variant of [18, Lemma 5]. Here, for any $C>0$, we set $\mathcal{X}_{C}=\left\{(\tau_{12},\dots,\tau_{34})\in\mathbb{R}^{6}_{\geq 0}:\,\sum_{1\leq i<j\leq 4}\tau_{ij}=C\right\}$.
Lemma 2.
Let $C>0$ and $\lambda\geq 1$. Then for each $(\tau_{12},\tau_{13},\tau_{14},\tau_{23},\tau_{24},0)\in\mathcal{X}_{C}$ we have
(4) $f(\lambda\tau_{12},\tau_{13},\tau_{14},\tau_{23},\tau_{24},0)\leq\frac{16C^{3}\lambda^{3}}{27(4\lambda-1)^{2}},$
with equality if and only if
$\tau_{12}=\frac{4\lambda-3}{2\lambda}\tau_{13}=\ldots=\frac{4\lambda-3}{2\lambda}\tau_{24}=\frac{2\lambda}{12\lambda-3}C.$
Proof.
Without loss of generality, we assume that $C=1$ and for simplicity, in the proof we denote the operator $\frac{\partial}{\partial\tau_{ij}}$ by $\partial_{ij}$. Given that $\mathcal{X}_{C}$ is compact, the function $f$ attains its maximum on it at some point $(\tau^{*},0)\in\mathcal{X}_{C}$ with $\tau^{*}=(\tau^{*}_{12},\ldots,\tau_{24}^{*})$.
Case 1: all coordinates of $\tau^{*}$ are positive.
Then, by the Lagrange multiplier method, the gradient of $f$ at $\tau^{*}$ is parallel to the gradient of the function $\sum_{i\neq j,ij\neq 34}\tau_{ij}-1$ defining the condition, thus, all five partial derivatives of $f$ at $\tau^{*}$ are equal. From the equations $(\partial_{13}f)(\tau^{*})=(\partial_{14}f)(\tau^{*})$ and $(\partial_{23}f)(\tau^{*})=(\partial_{24}f)(\tau^{*})$ and by the assumption that all coordinates of $\tau^{*}$ are positive, we obtain that $\tau_{13}^{*}=\tau_{14}^{*}$ and $\tau_{23}^{*}=\tau_{24}^{*}$. Eliminating $\tau_{14}^{*}$ and $\tau_{24}^{*}$ from the partial derivatives, the equation $(\partial_{13}f)(\tau^{*})=(\partial_{23}f)(\tau^{*})$, and the fact that all $\tau_{ij}^{*}$ as well as $\lambda$ are positive, yields that $\tau_{13}^{*}=\tau_{23}^{*}$. Thus, we need to find the maximum of $f(\lambda\tau_{12},\tau_{13},\ldots,\tau_{13},0)$ under the condition that $\tau_{12}+4\tau_{13}=1$. From this, by simple calculus methods, we obtain that $f(\lambda\tau_{12},\tau_{13},\ldots,\tau_{13},0)\leq\frac{16\lambda^{3}}{27(4\lambda-1)^{2}}$, with equality if and only if $\tau_{12}=\frac{4\lambda-3}{2\lambda}\tau_{13}=\frac{2\lambda}{12\lambda-3}$.
Case 2: $\tau_{12}^{*}=0$ and all other $\tau_{ij}^{*}$ are positive.
Note that $f(0,\tau_{1}3,\ldots,\tau_{24},0)=(\tau_{13}+\tau_{24})\tau_{14}\tau_{23}+(\tau_{14}+\tau_{23})\tau_{13}\tau_{24}$. In this case, by the same argument as in Case 1, we obtain that $\tau_{13}^{*}=\ldots=\tau_{24}^{*}=\frac{1}{4}$, and $f(\tau^{*})=\frac{1}{16}$.
Case 3: $\tau_{ij}^{*}=0$ for some $ij\neq 12$, and all other $\tau_{st}^{*}$ are positive.
Without loss of generality, we may assume that $\tau_{24}^{*}=0$. By a method similar to the one in Case 3, we obtain that if $0<\lambda\leq\frac{1}{2}$, there is no solution $\tau^{*}$ with positive coordinates, and if $\lambda>\frac{1}{2}$, then $\tau_{12}^{*}=\frac{2(2\lambda-1)}{3(4\lambda-1)}$, $\tau_{13}^{*}=\tau_{14}^{*}=\frac{2\lambda}{3(4\lambda-1)}$, and $\tau_{23}^{*}=0$, and $f(\tau^{*})=\frac{4\lambda^{2}}{27(4\lambda-1)}$.
Case 4, at least two $\tau_{ij}^{*}$ are zero.
We may assume that exactly two $\tau_{ij}^{*}$ are zero, and that there is no $s\in\{1,2,3,4\}$ such that the indices of all nonzero $\tau_{ij}^{*}$s contain $s$, as otherwise $f(\tau^{*},0)=0$. Under this condition, depending on which $\tau_{ij}^{*}$ are not zero, we have that $f(\tau^{*})=\frac{\lambda}{27}$ or $f(\tau^{*})=\frac{1}{27}$, with equality if and only if all nonzero $\tau_{ij}^{*}$ are equal to $\frac{1}{3}$.
As a summary, we need to find
$\max\left\{\frac{16\lambda^{3}}{27(4\lambda-1)^{2}},\frac{1}{16},\frac{4\lambda^{2}}{27(4\lambda-1)},\frac{\lambda}{27},\frac{1}{27}\right\}.$
Then an elementary computation shows that under the condition $\lambda\geq 1$, the maximum is $\frac{16\lambda^{3}}{27(4\lambda-1)^{2}}$, and the assertion follows. ∎
2. The proof of Theorem 1
Theorem 1 readily follows from Theorem 2 by Lemma 3.
Lemma 3.
Let $P$ be a unit volume parallelohedron in $\mathbb{R}^{3}$, and let $M$ be a tiling of $\mathbb{R}^{3}$ with translates of $P$. Set $m=(2,1)$. Then $\rho_{1}(\mathcal{M})\geq w_{m}(P)$, with equality if $\mathcal{M}$ is face-to-face.
Proof.
First, we assume that $M$ is face-to-face and that $P$ is a cell of $M$. Let us choose an edge $E$ of $P$. Let $\mathcal{E}$ contain the edges $E^{\prime}$ of $\mathcal{M}$ for which there is a sequence of cells $P_{1},\ldots,P_{k}$ such that
  1. (i)
    for any $i=1,2,\ldots,k-1$, $P_{i}\cap P_{i+1}$ is a face containing an edge that is a translate of $E$;
  2. (ii)
    $E$ is an edge of $P_{1}$ and $E^{\prime}$ is an edge of $P_{k}$.
Furthermore, let $\mathcal{P}$ be the family of cells containing edges from $\mathcal{E}$.
Note that the projection of the members of $\mathcal{P}$ in the direction of $E$ onto a plane $H$ is a translative, convex, edge-to-edge tiling of $H$. By the second property of parallelohedra in the list in the beginning of Section 1, the projections of the elements of $\mathcal{P}$ are either parallelograms, in which every vertex belongs to exactly four parallelograms, or they are centrally symmetric hexagons, in which every vertex belongs to three hexagons. Thus, a cell contains $4$ or $6$ translates of any given edge, and in the first case every edge belongs to exactly $4$ cells, and in the second case every edge belongs to exactly $3$ cells.
Now, we assign a weight $w_{E}$ to each edge $E$ of $M$; we set $w(E)=\frac{1}{k}$, where $k$ is the number of cells $E$ belongs to. Let $N_{r}$ denote the number of cells of $\mathcal{M}$, and $\mathcal{E}_{r}$ denote the family of the edges of the cells contained in $r\mathbf{B}^{3}$. As $N_{r}=\operatorname{vol}_{3}(r\mathbf{B}^{3})+\mathcal{O}(r^{2})$, we have
$L(\operatorname{skel}_{1}(\mathcal{M})\cap(r\mathbf{B}^{3}))=\sum_{E\in\mathcal{E}_{r}}L(E)+\mathcal{O}(r^{2})=N_{r}\sum_{i=1}^{6}w_{i}|p_{i}|+\mathcal{O}(r^{2}),$
implying Lemma 3 for face-to-face mosaics.
Finally, if $\mathcal{M}$ is a not necessarily face-to-face tiling by translates of $P$, then we may divide each edge into finitely many pieces in such a way that relative interior points on each piece belong to the same cells. It is easy to see that any part of an edge in a $6$-belt belongs to exactly $3$ cells, and a part of an edge in a $4$-belt belongs to either $4$ or $2$ cells. Thus, repeating the procedure, the weighted sum of the edge lengths in $r\mathbf{B}^{3}$, defined as in the previous case, yields a value for $\rho_{1}(M)$ asymptotically not less than that for a face-to-face mosaic generated by $P$. ∎
3. The proof of Theorem 2
First, for each value of $i$ and $m=(\alpha_{4},\alpha_{6})$, we find the infimum of $w_{m}(P)$ of the type (i) parallelohedra for each $i$ separately, which we denote by $w_{i}(\alpha_{4},\alpha_{6})$, then compare these values.
Case 1: $P$ is a type (1) parallelohedron.
Then $P$ is a parallelopiped, and it is easy to see that $w_{m}(P)$ is minimal if and only if $P$ is a unit cube, and in this case $w_{m}(P)=3\alpha_{4}$. Thus, $w_{1}(\alpha_{4},\alpha_{6})=3\alpha_{4}$.
Case 2: $P$ is a type (2) parallelohedron.
Then $P$ is a centrally symmetric hexagon-based prism. Clearly, we can assume that $P$ is a right prism, and by the Discrete Isoperimetric Inequality, we may assume that its base is a regular hexagon. Let the edge length of the base be $a$, and the length of the lateral edges be $b$. Then $\operatorname{vol}_{3}(P)=\frac{3\sqrt{3}}{2}a^{2}b$, and $w_{m}(P)=3\alpha_{4}a+\alpha_{6}b$. An elementary computation shows that if $\frac{3\sqrt{3}}{2}a^{2}b=1$, then $w_{m}(P)\geq\frac{3^{7/6}\alpha_{4}^{2/3}\alpha_{6}^{1/3}}{2^{1/3}}$, with equality if and only if $a=\frac{2^{2/3}\alpha_{6}^{1/3}}{3^{5/6}\alpha_{4}^{1/3}}$ and $b=\frac{3^{1/6}\alpha_{4}^{2/3}}{2^{1/3}\alpha_{6}^{2/3}}$. Hence, $w_{2}(\alpha_{4},\alpha_{6})=\frac{3^{7/6}\alpha_{4}^{2/3}\alpha_{6}^{1/3}}{2^{1/3}}$.
Case 3: $P$ is a type (3) parallelohedron.
In this case $P$ is generated by $4$ segments, and each segment generates a $6$-belt. Thus, $w_{m}(P)=\frac{\alpha_{6}}{2}w(P)$, where $w(P)$ denotes the mean width of $P$. This means that we may apply the proof of [18, Theorem 1] (see also Table 1 of [18]), and obtain that $w_{m}(P)\geq 3^{1/2}2^{2/3}\alpha_{6}$, with equality if and only if $P$ is a regular rhombic dodecahedron, implying, in particular, that $w_{3}(\alpha_{4},\alpha_{6})=3^{1/2}2^{2/3}\alpha_{6}$.
Case 4: $P$ is a type (4) parallelohedron.
First, note that if $\alpha_{4}>\alpha_{6}$, then $w_{4}(\alpha_{4},\alpha_{6})\geq w_{4}(\alpha_{6},\alpha_{6},)>w_{5}(\alpha_{6},\alpha_{6})$, where the last inequality follows from [18, Theorem 1]. In the following we assume that $\alpha_{4}\leq\alpha_{6}$.
To prove Theorem 2 in this case, we use the notation introduced in Section 1: we represent $P$ in the form
$P=\sum_{1\leq i<j\leq 4}[o,\beta_{ij}v_{i}\times v_{j}]$
for some $v_{1},v_{2},v_{3},v_{4}\in\mathbb{R}^{3}$ satisfying $\sum_{i=1}^{4}v_{i}=o$, and for some $\beta_{ij}\geq 0$, with exactly one $\beta_{ij}$ equal to zero. By our assumptions, $|V_{ijk}|=1$ for any $\{i,j,k\}\subset\{1,2,3,4\}$, where $V_{ijk}$ denotes the determinant with $v_{i},v_{j},v_{k}$ as columns, and we assumed that $V_{123}=1$, which implies, in particular, that $V_{124}=-1$. Then $T=\operatorname{conv}\{v_{1},v_{2},v_{3},v_{4}\}$ is a centered tetrahedron, with volume $\operatorname{vol}_{3}(T)=\frac{2}{3}$.
Without loss of generality, we set $\beta_{34}=0$. Then, among the edges of $P$, the translates of $\beta_{12}v_{1}\times v_{2}$ belong to a $4$-belt, and the translates of all other edges belong to $6$-belts. Thus, we have
$w_{m}(P)=\alpha_{4}\beta_{12}|v_{1}\times v_{2}|+\alpha_{6}\sum_{i\in\{1,2\},j\in\{3,4\}}\beta_{ij}|v_{i}\times v_{j}|,\hbox{ and}$
$\operatorname{vol}_{3}(P)=f(\beta_{12},\beta_{13},\ldots,\beta_{24},0).$
We follow the two steps of the proof of [18, Theorem 1]. Our first step is to find the minimum of $w_{m}(P)$ in the affine class of $P$. For this, we use Lemma 4, proved by Petty [22] (see also (3.4) in [11] for arbitrary convex bodies).
Lemma 4.
Let $P\subset\mathbb{R}^{n}$ be a convex polytope with outer unit facet normals $u_{1},\ldots,u_{k}$. Let $F_{i}$ denote the $(n-1)$-dimensional volume of the $i$th facet of $P$. Then, up to congruence, there is a unique volume preserving affine transformation $A$ such that the surface area $\operatorname{surf}(A(P))$ is maximal in the affine class of $P$. Furthermore, $P$ satisfies this property if and only if its surface area measure is isotropic, that is, if
(5) $\sum_{i=1}^{k}\frac{nF_{i}}{\operatorname{surf}(P)}u_{i}\otimes u_{i}=\operatorname{Id}$
where $\operatorname{Id}$ denotes the identity matrix.
Any convex polytope satisfying the conditions in (5) is said to be in surface isotropic position. We note that the volume of the projection body of any convex polyhedron is invariant under volume preserving linear transformations (cf. [23]). On the other hand, from Cauchy’s projection formula and the additivity of the support function (see [10]) it follows that the projection body of the polytope in Lemma 4 is the zonotope $\sum_{i_{1}}^{k}[o,F_{i}u_{i}]$. Note that the solution to Minkowski’s problem [24] states that some unit vectors $u_{1},\ldots,u_{k}\in\mathbb{R}^{n}$ and positive numbers $F_{1},\ldots,F_{k}$ are the outer unit normals and volumes of the facets of a convex polytope if and only if the $u_{i}$s span $\mathbb{R}^{n}$, and $\sum_{i=1}^{k}F_{i}u_{i}=o$. This yields that there is an $o$-symmetric convex polyhedron $Q$ whose faces have outer unit normals ±$\pm\frac{v_{1}\times v_{2}}{|v_{1}\times v_{2}|}$, and ±$\pm\frac{v_{i}\times v_{j}}{|v_{i}\times v_{j}|}$ with $ij=13,23,14,24$, and these faces have area $\frac{1}{2}\alpha_{4}\beta_{12}|v_{1}\times v_{2}|$ and $\frac{1}{2}\alpha_{6}\beta_{ij}|v_{i}\times v_{j}|$ with $ij=13,14,23,24$, respectively. Since $u\otimes u=(-u)\otimes(-u)$ for any $u\in\mathbb{R}^{n}$, we may apply Lemma 4 for $Q$, and obtain that
(6) $\frac{\alpha_{4}\beta_{12}}{|v_{1}\times v_{2}|}(v_{1}\times v_{2})\otimes(v_{1}\times v_{2})+\alpha_{6}\sum_{i\in\{1,2\},j\in\{3,4\}}\frac{2\beta_{ij}}{|v_{i}\times v_{j}|}(v_{i}\times v_{j})\otimes(v_{i}\times v_{j})=\frac{w_{m}(P)}{3}\operatorname{Id},$
which in the following we assume to hold for $P$.
Recall that $x\otimes y=xy^{T}$ and $\langle x,y\rangle=x^{T}y$ for any column vectors $x,y\in\mathbb{R}^{n}$. Hence, multiplying both sides in (6) by $v_{3}^{T}$ from the left and $v_{4}$ from the right, it follows that
$\langle v_{3},v_{4}\rangle=\frac{3\alpha_{4}\beta_{12}}{w_{m}(P)|v_{1}\times v_{2}|}V_{123}V_{124}.$
Since $V_{123}=-V_{124}=1$, we have
$\beta_{12}=-\langle v_{3},v_{4}\rangle|v_{1}\times v_{2}|\cdot\frac{w_{m}(P)}{3\alpha_{4}}.$
Multiplying both sides in (6) with some other $v_{i}$s, we obtain that for any $i\in\{1,2\}$ and $j\in\{3,4\}$, and with the notation $\{i,j,s,t\}=\{1,2,3,4\}$, we have
$\beta_{ij}=-\langle v_{s},v_{t}\rangle|v_{i}\times v_{j}|\cdot\frac{w_{m}(P)}{3\alpha_{6}},$
and obtain similarly that $\langle v_{1},v_{2}\rangle=0$.
Now we set $\zeta_{ij}=-\langle v_{s},v_{t}\rangle|v_{i}\times v_{j}|$ for all $\{i,j,s,t\}=\{1,2,3,4\}$, and note that $\beta_{12}=\frac{w_{m}(P)\zeta_{12}}{3\alpha_{4}}$, $\beta_{34}=\zeta_{34}=0$, and for any $ij\neq 12,34$, $\beta_{ij}=\frac{w_{m}(P)\zeta_{ij}}{3\alpha_{6}}$. Substituting back these quantities into the formulas for $\operatorname{vol}_{3}(P)$ and $w_{m}(P)$ and simplifying, we may rewrite our optimization problem in the following form: find the maximum of $\frac{27\operatorname{vol}_{3}(P)}{\left(w_{m}(P)\right)^{3}}=f\left(\frac{\zeta_{12}}{\alpha_{4}},\frac{\zeta_{13}}{\alpha_{6}},\ldots,\frac{\zeta_{34}}{\alpha_{6}},0\right)$, where $\zeta_{ij}$ is defined as above, in the family of all centered tetrahedra $T$ with $\operatorname{vol}_{3}(T)=\frac{2}{3}$, under the condition that $\sum_{1\leq i<j\leq 4}\zeta_{ij}|v_{i}\times v_{j}|=3$. Here, it is worth noting that the last condition is satisfied for any centered tetrahedron $T$ with volume $\frac{2}{3}$ by Lemma 1, and thus, it is redundant. Note that since $f$ is $3$-homogeneous, we have
$f\left(\frac{\zeta_{12}}{\alpha_{4}},\frac{\zeta_{13}}{\alpha_{6}}\ldots,\frac{\zeta_{34}}{\alpha_{4}},0\right)=\frac{1}{\alpha_{6}^{3}}f\left(\mu\zeta_{12},\zeta_{13},\ldots,\zeta_{34},0\right),$
where $\mu=\frac{\alpha_{6}}{\alpha_{4}}\geq 1$.
In the second step we give an upper bound for the values of this function. Set $\gamma_{ij}=-\langle v_{s},v_{t}\rangle$ and $\tau_{ij}=\gamma_{ij}|v_{i}\times v_{j}|^{2}$ for all $\{i,j,s,t\}=\{1,2,3,4\}$, and note that this yields $\gamma_{34}=\tau_{34}=0$. To find an upper bound, we apply the Cauchy-Schwartz Inequality, which states that for any real numbers $x_{i},y_{i}$, $i=1,2,\ldots,k$, we have $\sum_{i=1}^{k}x_{i}y_{i}\leq\sqrt{\sum_{i=1}^{k}x_{i}^{2}}\sqrt{\sum_{i=1}^{k}y_{i}^{2}}$, with equality if and only if the $x_{i}$s and the $y_{i}$s are proportional. To do this, we write each member of $f$ not containing $\zeta_{12}$ as the product $\zeta_{ij}\zeta_{kl}\zeta_{mn}=\sqrt{\gamma_{ij}\gamma_{kl}\gamma_{mn}}\sqrt{\tau_{ij}\tau_{kl}\tau_{mn}}$, and each member containing $\zeta_{12}$ as $\mu\zeta_{12}\zeta_{kl}\zeta_{mn}=\sqrt{\gamma_{12}\gamma_{kl}\gamma_{mn}}\sqrt{\mu^{2}\tau_{12}\tau_{kl}\tau_{mn}}$. Thus, we obtain that
$f(\mu\zeta_{12},\zeta_{13},\ldots,\zeta_{24},0)\leq\sqrt{f(\mu^{2}\tau_{12},\tau_{13},\tau_{14},\tau_{23},\tau_{24},0)}\cdot\sqrt{f(\gamma_{12},\gamma_{13},\gamma_{14},\gamma_{23},\gamma_{24},\gamma_{34})}.$
Here we use the fact that by Lemma 1, $f(\gamma_{12},\ldots,\gamma_{34})=1$ for all centered tetrahedra with volume $\frac{2}{3}$. Furthermore, observe that by Lemma 1 we have $\sum_{1\leq i\leq j\leq 4}\tau_{ij}=3$ for all such tetrahedra. Hence, by Lemma 2, we have
$f(\mu^{2}\tau_{12},\tau_{13},\ldots,\tau_{24},0)\leq\frac{16\mu^{6}}{(4\mu^{2}-1)^{2}},$
implying that if $\operatorname{vol}_{3}(P)=1$, then
$w_{m}(P)\geq\frac{3\alpha_{4}^{1/3}(4\alpha_{6}^{2}-\alpha_{4}^{2})^{1/3}}{2^{2/3}}.$
Consequently, we have
(7) $w_{4}(\alpha_{4},\alpha_{6})\geq\left\{\begin{array}[]{l}w_{5}(\alpha_{6},\alpha_{6}),\hbox{ if }\alpha_{4}>\alpha_{6};\\ \frac{3\alpha_{4}^{1/3}(4\alpha_{6}^{2}-\alpha_{4}^{2})^{1/3}}{2^{2/3}},\hbox{if }\alpha_{4}\leq\alpha_{6},\end{array}\right.$
and equality in the first case does not occur.
Case 5: $P$ is a type (5) parallelohedron. In this case every generating segment generates a $6$-belt, and we can apply [18, Theorem 1] directly, and obtain that as $\operatorname{vol}_{3}(P)=1$,
$w_{m}(P)\geq\frac{3}{2^{1/6}}\alpha_{6},$
with equality if and only if $P$ is a regular truncated octahedron.
Table 1 collects our lower bounds for $w_{m}(P)$ for the different types of parallelohedra. Note that if $i\neq 4$, $w_{i}(\alpha_{4},\alpha_{6})$ coincides with the expression in the $i$th row and second column of Table 1, whereas $w_{4}(\alpha_{4},\alpha_{6})$ is not less than the quantity in the $4$h row and the second column.
Type
$w_{i}(\alpha_{4},\alpha_{6})$
Optimal parallelohedra
$3\alpha_{4}$
unit cube
$\frac{3^{7/6}}{2^{1/3}}\alpha_{4}^{2/3}\alpha_{6}^{1/3}$
regular hexagon based right prism, with base and lateral edges of lengths $\frac{2^{2/3}\alpha_{6}^{1/3}}{3^{5/6}\alpha_{4}^{1/3}}$ and $\frac{3^{1/6}\alpha_{4}^{2/3}}{2^{1/3}\alpha_{6}^{2/3}}$, respectively
$2^{2/3}\cdot 3^{1/2}\alpha_{6}$
regular rhombic dodecahedron with edge length $\frac{\sqrt{3}}{2^{4/3}}$
$\geq\frac{3\alpha_{4}^{1/3}(4\alpha_{6}^{2}-\alpha_{4}^{2})^{1/3}}{2^{2/3}}$, or $>w_{5}(\alpha_{6},\alpha_{6})$
not known
$\frac{3}{2^{1/6}}\alpha_{6}$
regular truncated octahedron of edge length $\frac{1}{2^{7/6}}$
Table 1. Minima of the weighted mean widths of different types of parallelohedra
In the final step we determine the minimum of $\{w_{i}(\alpha_{4},\alpha_{6}):i=1,2,3,4,5\}$ as a function of $\alpha_{4}$ and $\alpha_{6}$.
Figure 2. The lower bounds for $w_{i}(\alpha_{4},\alpha_{6})$ from the second column of Table 1, with $\alpha_{6}=1$, and $0<\alpha_{4}\leq\alpha_{6}$.
An elementary computation shows that $w_{3}(\alpha_{4},\alpha_{6})>w_{5}(\alpha_{4},\alpha_{6})$ for all $\alpha_{4},\alpha_{6}>0$, and that $w_{4}(\alpha_{4},\alpha_{6})\geq\frac{3\alpha_{4}^{1/3}(4\alpha_{6}^{2}-\alpha_{4}^{2})^{1/3}}{2^{2/3}}>\min\{w_{1}(\alpha_{4},\alpha_{6}),w_{5}(\alpha_{4},\alpha_{6})\}$ if $0<\alpha_{4}\leq\alpha_{6}$. Furthermore, if $\alpha_{4}>\alpha_{6}$, then $w_{4}(\alpha_{4},\alpha_{6})>w_{5}(\alpha_{6},\alpha_{6})=w_{5}(\alpha_{4},\alpha_{6})$. Thus, we need to compare only the values of $w_{1}(\alpha_{4},\alpha_{6})$, $w_{2}(\alpha_{4},\alpha_{6})$ and $w_{5}(\alpha_{4},\alpha_{6})$, and from this another elementary computation proves the statement.
4. Proof of Theorem 3
Let $\mathcal{M}\in\mathcal{F}_{dc}$. Then $\mathcal{M}$ can be decomposed into a planar mosaic $\mathcal{M}_{1}$ in a plane $H$ and a linear mosaic $\mathcal{M}_{2}$ in a line $L$. Since every cell in $\mathcal{M}$ has volume one, $L$ is decomposed into congruent segments, and $H$ is tiled with equal area convex polygons. To minimize the edge density of $\mathcal{M}$, we may assume that $L$ is orthogonal to $H$. Let the length of the cells in $\mathcal{M}_{2}$ be denoted by $\ell$. Since $\mathcal{M}$ has unit volume cells, the area of the cells of $\mathcal{M}_{1}$ is $\frac{1}{\ell}$. Since $\mathcal{M}$ is normal, it is easy to see that $\mathcal{M}_{1}$ is normal. We permit the cells of $\mathcal{M}_{1}$ to have straight angles, and in this way we can regard $\mathcal{M}_{1}$ as an edge-to-edge mosaic.
In the following we assume that $H$ coincides with the $(x,y)$-coordinate plane, and $L$ coincides with the $z$-axis. Note that by the symmetry of the mosaic, the edge density of $\mathcal{M}$ is equal to the edge density of the part of $\mathcal{M}$ in the infinite, half open strip $C=\{0\leq z<m\}$.
Consider the half-open right cylinder $C(R)=\{(x,y,z)\in\mathbb{R}^{3}:x^{2}+y^{2}\leq R^{2},0\leq z<\ell\}$ of volume $R^{2}\pi\ell$. We estimate the total length of the edges of $\mathcal{M}$ in $C(R)$. Let $k$ and $N$ denote the numbers of cells and nodes of $\mathcal{M}_{1}$ in $C(R)$, respectively. Let the cells of $\mathcal{M}_{0}$ in $C(R)$ be $F_{1},F_{2},\ldots,F_{k}$, and let $v_{i}$ be the number of the edges of $F_{i}$. Let $S$ denote the sum of all interior angles of all cells of $\mathcal{M}_{1}$ in $C(R)\cap H$. Then, on one hand,
$S=2\pi N+\mathcal{O}(R),$
and on the other hand, it is
$\sum_{i=1}^{k}(v_{i}-2)\pi+\mathcal{O}(R).$
Since $\operatorname{area}(F_{i})=\frac{1}{\ell}$ for all $i$, $k=R^{2}\pi\ell+\mathcal{O}(R)$. This yields that
$N=\frac{R^{2}\pi\ell}{2}\left(\frac{\sum_{i=1}^{k}v_{i}}{k}-2\right)+\mathcal{O}(R).$
Thus, the total edge length of $\mathcal{M}$ in $C(R)$ is
$L(\operatorname{skel}_{1}(\mathcal{M}_{1})\cap C(R))=\ell N+\frac{1}{2}\sum_{i=1}^{k}\operatorname{perim}(F_{i})+\mathcal{O}(R),$
where $\operatorname{perim}(\cdot)$ denotes perimeter. Since $\operatorname{vol}_{3}(C(R))=R^{2}\pi\ell$, we have that
$\rho_{1}(\mathcal{M})=\frac{\ell}{2}(v-2)+\frac{1}{2}\lim_{R\to\infty}\frac{\sum_{i=1}^{k}\operatorname{perim}(F_{i})}{k},$
where $v=\lim_{R\to\infty}\frac{\sum_{i=1}^{k}}{k}$ is the average number of edges of a cell of $\mathcal{M}$, which exists, since $\mathcal{M}$ is normal.
By the Discrete Isoperimetric Inequality, $\operatorname{perim}(F_{i})\geq 2\sqrt{v_{i}\tan\frac{\pi}{v_{i}}}\cdot\frac{1}{\sqrt{\ell}}$; here the right-hand side is the perimeter of the regular $v_{i}$-gon of area $\frac{1}{\ell}$. By the convexity of the function $x\mapsto\sqrt{x\tan\frac{\pi}{x}}$ on the interval $[3,\infty)$, we have
(8) $\rho_{1}(\mathcal{M})=\frac{\ell}{2}(v-2)+\sqrt{v\tan\frac{\pi}{v}}\cdot\frac{1}{\sqrt{\ell}}.$
For any $3\leq v\leq 6$, the expression on the right-hand side of (8) is minimal if $\ell=\frac{\sqrt[3]{v\tan\frac{\pi}{v}}}{(v-2)^{2/3}}$. Substituting it back into (8), we obtain that
$\rho_{1}(\mathcal{M})\geq\frac{3}{2}\left(v(v-2)\tan\frac{\pi}{v}\right)^{1/3}.$
Then the assertion follows from the facts that the expression on the right-hand side is strictly increasing on the interval $[3,6]$, and that an average cell in a planar mosaic has at most $6$ edges.